OpenStudy (anonymous):

a^x - b^x / x as x tends to 0

OpenStudy (jamesj):

Do you know the derivative of a^x wrt (with respect to) x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

xlna?

OpenStudy (anonymous):

would we apply l'hospital rule?

OpenStudy (jamesj):

Good then write \[\frac{a^x - b^x}{x} = \frac{a^x - 1}{x - 0} - \frac{b^x - 1}{x - 0}\] Then notice the the limit of each of these terms is the derivate of a^x and b^x respectively evaluated at x=0.

OpenStudy (anonymous):

yea now it is simple

OpenStudy (anonymous):

thank u

OpenStudy (jamesj):

Sure thing.